Derived Newton’s second law using constant acceleration ( Distance of Free-Fall) and find Work is equivalent to KE.
Answer #1

according to newton's second law

force applied is equal to the rate of change of momentum

F = dp/dt

F = mdv/dt

dv/dt = a

F = ma -------(1)

dW = F*dS

dW = ma *dS

dW = mv(dv/ds)*ds

dW = mv dv

so integrating both sides we have

W = 0.5m(v2^2 - v1^2)

so work done = change in Kinetic energy